OpenStudy (anonymous):

could someone tell me if I'm doing this problem correctly Given the function f(x) = 4(2)x, Section A is from x = 1 to x = 2 and Section B is from x = 3 to x = 4. Part A: Find the average rate of change of each section. Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other.

OpenStudy (anonymous):

please

jimthompson5910 (jim_thompson5910):

Can you post your work so far?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

part a f(x)= 4(2)^2-4(2)^1 =8

jimthompson5910 (jim_thompson5910):

you will then divide that by the change in x, which is 2-1 = 1

OpenStudy (anonymous):

Part B 4(2)^4-4(2)^3=32

OpenStudy (anonymous):

ok so i just put that underneath like a fraction so would it be 8 over 1

jimthompson5910 (jim_thompson5910):

yes, because the avg rate of change is equal to the change in y over the change in x

OpenStudy (anonymous):

oh ok so i have to write it 8 over 1 same thing with b 32over 1

jimthompson5910 (jim_thompson5910):

since 8/1 = 8, the final answer to section A is simply 8

OpenStudy (anonymous):

oh ok but im having a hard time putting it in words to explain why b is greater than a in part B where it says explain

jimthompson5910 (jim_thompson5910):

what does the graph of 4*(2)^x look like?

OpenStudy (anonymous):

 I cant draw it very good on here